OpenStudy (anonymous):

Solve the equation. Type just the numeral answer. Do NOT round your solution.n/-1.5=-1.2 n=

OpenStudy (anonymous):

1.8

OpenStudy (anonymous):

i surebecuz i gt 1.25

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

n/1.5=-1.2 (-2/3)n=-1,2 divide both by -2/3 leads to n=-1,2/-2/3 n=1.8

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

-2/3 is the invers btw